cos a + cos b – cos (a+b) = 3/2 show that a=b= 60 degrees

Grade:11

1 Answers

Ahsan Barkati
9 years ago
Let A+B+C=180
 cosA+cosB+cosC = cos A+ cos B +cos A+B
cosA+cos B- cosC= cos A+ cos B -cos A+B
                            3/2= 1-4sinA/2 sin B/2 cosC/2
4sinA/2 sin B/2 cosA+B/2=--1/2
A=b=60